# Geodesic Synthetic Control Methods for Random Objects and Functional Data

### v1

@@ 6.1. Construction of the estimator
We discuss an extension of synthetic difference-in-differences developed in Arkhangelsky et al., (2021) for the Euclidean case to outcomes located in geodesic metric spaces. Define
$Y_{j,T_{0}+1:T}^{(\oplus)}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{j,t}),\quad j=1,\dots,J+1,$
$Y_{1,T_{0}+1:T}^{(\oplus)}(D)$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{1,t}(D)),\quad D\in\{N,I\}.$
@@
$Y_{2:J+1,t}^{(\bm{w})}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t}),\quad Y_{j,1:T_{0}}^{(\bm{\lambda})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,Y_{j,t}),$
$Y_{2:J+1,1:T_{0}}^{(\bm{w},\bm{\lambda})}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,1:T_{0}}^{(\bm{\lambda})}),\quad Y_{2:J+1,T_{0}+1:T}^{(\bm{w},\oplus)}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,T_{0}+1:T}^{(\oplus)}).$
With $\Gamma_{\alpha,\beta}:\mathcal{M}\to\mathcal{M}$ denoting a geodesic transport map for $\alpha,\beta\in\mathcal{M}$ as introduced in Assumption 3.2, we define the GSDID estimator for $\tau$ as
@@
$\overline{\bm{w}}\in\mathop{\rm arg~{}min}\limits_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})}),\quad\overline{\bm{\lambda}}\in\mathop{\rm arg~{}min}\limits_{\bm{\lambda}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,1:T_{0}}^{(\bm{\lambda})}).$
@@
which is the SDID estimator of Arkhangelsky et al., (2021). Thus, the proposed estimator is a natural extension of the established case with Euclidean outcomes. We discuss extensions to cases with multiple treated units or where the interest lies in the causal effects at multiple post-treatment time points in Remark 6.3 below.
Remark 6.1 (Connection with geodesic synthetic control method and geodesic difference-in-differences).
To discuss the relation of the proposed GSDID estimator with the GSC estimator introduced above and the geodesic DID (GDID) estimator introduced in Zhou et al., (2025), we note that GSC, GDID, and GSDID estimators can be represented as geodesic transports based on weighted averages of outcomes $\delta_{j,k}^{(\mathrm{M})}$, $k=1,2,3$, using appropriate weights $w_{j}^{(\mathrm{M})}$, $\mathrm{M}\in\{\mathrm{GSC},\mathrm{GDID},\mathrm{GSDID}\}$, i.e.,
@@
$Y_{1,T_{0}+1:T}^{(\mathrm{M})}=\Gamma_{\delta_{2:J+1,1:T_{0}}^{(\mathrm{M})},\delta_{2:J+1,T_{0}+1:T}^{(\mathrm{M})}}\left(\delta_{1,1:T_{0}}^{(\mathrm{M})}\right),\quad\delta_{2:J+1,1:T_{0}}^{(\mathrm{M})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}^{(\mathrm{M})}d^{2}(\nu,\delta_{j,1:T_{0}}^{(\mathrm{M})}),$
$\delta_{2:J+1,T_{0}+1:T}^{(\mathrm{M})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}^{(\mathrm{M})}d^{2}(\nu,\delta_{j,T_{0}+1:T}^{(\mathrm{M})}),\quad\delta_{1,1:T_{0}}^{(\mathrm{M})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}^{(\mathrm{M})}d^{2}(\nu,\overline{\delta}_{j,1:T_{0}}^{(\mathrm{M})}),$
@@
$\{w_{j}^{(\mathrm{GSC})},\delta_{j,1:T_{0}}^{(\mathrm{GSC})},\delta_{j,T_{0}+1:T}^{(\mathrm{GSC})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSC})}\}$ =$=\{\overline{w}_{j},Y_{j,1:T_{0}}^{(\oplus)},Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,1:T_{0}}^{(\oplus)}\},$
$\{w_{j}^{(\mathrm{GDID})},\delta_{j,1:T_{0}}^{(\mathrm{GDID})},\delta_{j,T_{0}+1:T}^{(\mathrm{GDID})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GDID})}\}$ =$=\{\frac{1}{J},Y_{j,1:T_{0}}^{(\oplus)},Y_{j,T_{0}+1:T}^{(\oplus)},Y_{1,1:T_{0}}^{(\oplus)}\},$
$\{w_{j}^{(\mathrm{GSDID})},\delta_{j,1:T_{0}}^{(\mathrm{GSDID})},\delta_{j,T_{0}+1:T}^{(\mathrm{GSDID})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSDID})}\}$ =$=\{\overline{w}_{j},Y_{j,1:T_{0}}^{(\overline{\bm{\lambda}})},Y_{j,T_{0}+1:T}^{(\oplus)},Y_{1,1:T_{0}}^{(\overline{\bm{\lambda}})}\}.$
@@
$Y_{1,T_{0}+1:T}^{(\mathrm{GSC})}$ $=\Gamma_{Y_{2:J+1,1:T_{0}}^{(\overline{\bm{w}},\oplus)},Y_{2:J+1,T_{0}+1}^{(\overline{\bm{w}})}}\left(Y_{2:J+1,1:T_{0}}^{(\overline{\bm{w}},\oplus)}\right)=Y_{2:J+1,T_{0}+1}^{(\overline{\bm{w}})},$ (6.1)
$Y_{1,T_{0}+1:T}^{(\mathrm{GDID})}$ $=\Gamma_{Y_{2:J+1,1}^{(\oplus)},Y_{2:J+1,2}^{(\oplus)}}\left(Y_{1,1}\right),$ (6.2)
@@
$Y_{2:J+1,t}^{(\oplus)}$ $=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(\nu,Y_{j,t}),\ t=1,2.$
The resulting estimators (6.1) and (6.2) coincide with the GSC and GDID estimators, respectively. Under Assumption 6.2 below, we have $Y_{1,T_{0}+1}^{(\mathrm{GSC})}=Y_{1,T_{0}+1}(N)$. Comparing $\{\delta_{j,1:T_{0}}^{(\mathrm{GSC})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSC})}\}$ and $\{\delta_{j,1:T_{0}}^{(\mathrm{GSDID})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSDID})}\}$, the GSDID can be interpreted as attenuating the bias of the GSC estimator by introducing additional flexibility through the weights $\overline{\bm{\lambda}}$ that are not present in the GSC. Under Assumption 6.3 below, we have $Y_{1,2}^{(\mathrm{GDID})}=Y_{1,2}(N)$. Comparing the weights $w_{j}^{(\mathrm{GDID})}$ and $w_{j}^{(\mathrm{GSDID})}$, GSDID can be seen as introducing additional flexibility to estimate the target $Y_{1,T_{0}+1:T}^{(\oplus)}(N)$ through the weights $\overline{\bm{w}}$, which may attenuate bias when the parallelity assumption does not hold. The proposed GSDID thus emerges as a method that combines favorable properties of both GSC and GDID.
6.2. Causal model for geodesic synthetic difference-in-differences
@@
$Y_{j,t}(N)=g(U_{j},V_{t}),\quad t=1,\ldots,T,\ j=1,\ldots,J+1,$ (6.3)
@@
$g^{(\oplus)}(U_{j},V_{T_{0}+1:T})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}\!\!\!d^{2}(\nu,g(U_{j},V_{t})).$
@@
$g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,g(U_{j},V_{t})),$
$g^{(\bm{w})}(U_{2:J+1},V_{t})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g(U_{j},V_{t})),$
$g^{(\bm{w},\oplus)}(U_{2:J+1},V_{T_{0}+1:T})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g^{(\oplus)}(U_{j},V_{T_{0}+1:T})),$
$g^{(\bm{w},\bm{\lambda})}(U_{2:J+1},V_{1:T_{0}})=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})).$
@@
Assumption 6.1.
For any $\bm{w}\in\Delta^{J-1}$ and $\bm{\lambda}\in\Delta^{T_{0}-1}$, the objects $g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})$, $g^{(\oplus)}(U_{j},V_{T_{0}+1:T})$, $g^{(\bm{w})}(U_{2:J+1},V_{j})$, $g^{(\bm{w},\oplus)}(U_{2:J+1},V_{T_{0}+1:T})$, and $g^{(\bm{w},\bm{\lambda})}(U_{2:J+1},V_{1:T_{0}})$ exist and are unique.
@@
$\mathbb{W}_{t}^{*}$ =$=\!\{\bm{w}^{*}\!\in\!\Delta^{J-1}\!:\,\,\!d(g(U_{1},V_{t}),g^{(\bm{w}^{*})}(U_{2:J+1},V_{t}))\!=\!\inf_{\bm{w}\in\Delta^{J-1}}\!\!\!d(g(U_{1},V_{t}),g^{(\bm{w})}(U_{2:J+1},V_{t}))\},$
$\Lambda_{j}^{*}$ =$=\!\left\{\!\bm{\lambda}^{*}\!\in\!\Delta^{T_{0}-1}\!:\,\!d(g^{(\oplus)}(U_{j},V_{T_{0}+1:T}),g^{(\bm{\lambda}^{*})}(U_{j},V_{1:T_{0}}))\!=\!\inf_{\bm{\lambda}\in\Delta^{T_{0}-1}}\!\!\!\!d(g^{(\oplus)}(U_{j},V_{T_{0}+1:T}),g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}}))\!\right\}.$
Assumption 6.2 (Synthetic control).
@@
    $\mathbb{W}_{\bm{\lambda}}^{*}$ =$=\left\{\bm{w}\in\Delta^{J-1}:d(g^{(\bm{\lambda})}(U_{1},V_{1:T_{0}}),g^{(\bm{w},\bm{\lambda})}(U_{2:J+1},V_{1:T_{0}}))=0\right\},$
    $\mathbb{W}_{\oplus}^{*}$ =$=\left\{\bm{w}\in\Delta^{J-1}:d(g^{(\oplus)}(U_{1},V_{T_{0}+1:T}),g^{(\bm{w},\oplus)}(U_{2:J+1},V_{T_{0}+1:T}))=0\right\}.$
Assumption 6.3 (Parallel trend).
@@
Assumption 6.2 is for the GSC. Assumption 6.2 (ii) implies that the optimal weights $\bm{w}$ such that $Y_{2:J+1,t}^{(\bm{w})}=Y_{1,t}(N)$ at each time point $t=1,\dots,T_{0}$ are still optimal in the sense that $Y_{1,1:T_{0}}^{(\bm{\lambda})}(N)$ and $Y_{1,T_{0}+1:T}^{(\oplus)}(N)$ respectively coincide with $Y_{2:J+1,1:T_{0}}^{(\bm{w},\bm{\lambda})}(N)$ and $Y_{2:J+1,T_{0}+1:T}^{(\bm{w},\oplus)}(N)$ as long as $\bm{\lambda}$ belongs to the set of optimal weights $\bigcap_{j=2}^{J+1}\Lambda_{j}^{*}$ such that for $j=2,\dots,J+1$, $Y_{j,1:T_{0}}^{(\bm{\lambda})}$ is the closest to $Y_{j,T_{0}+1:T}^{(\oplus)}(N)$ with respect to the metric $d$ among the Fréchet means $Y_{j,1:T_{0}}^{(\bm{\lambda}_{0})}$ with $\bm{\lambda}_{0}\in\Delta^{T_{0}-1}$. Figure 7 illustrates the GSDID under Assumption 6.2.
Assumption 6.3 (ii) corresponds to the parallel trend assumption for the GDID. We refer to Zhou et al., (2025) for more details about its construction and theoretical properties. For Euclidean outcomes, Assumption 6.3 (ii) reduces to
@@
which corresponds to the standard assumption of parallel trends in the canonical DID framework (Zhou et al.,, 2025). Figure 8 illustrates the GSDID under Assumption 6.3.
For the space of networks (Example 2.1), Wasserstein space (Example 2.4), the space of functional data (Example 2.5), and the space of SPD matrices (Example 2.3) with either the Frobenius metric, the power metric family, the Log-Euclidean metric or the Log-Cholesky metric, one can verify Assumptions 6.1, 6.2, and 6.3 (i). Indeed, letting $g(U_{j},V_{t})=\gamma_{U_{j},V_{t}}(\alpha)$, $\alpha\in(0,1)$,
@@
$U_{2:J+1}^{(\bm{w})}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,U_{j}),$
$V_{T_{0}+1:T}^{(\oplus)}=\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}\!\!\!d^{2}(\nu,V_{t}),\ V_{1:T_{0}}^{(\bm{\lambda})}\in\mathop{\rm arg~{}min}\limits_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,V_{t}).$
For the space of compositional data (Example 2.2), consider the causal model $g(U_{j},V_{t})=\mathrm{Exp}_{V_{t}}(\mathrm{Log}_{U_{1}}(U_{j}))$. If we additionally assume that

### v2

@@ 7 Data Availability Statement
We discuss an extension of synthetic difference-in-differences developed in Arkhangelsky et al. (2021) for the Euclidean case to outcomes located in geodesic metric spaces. Define
$Y_{j,T_{0}+1:T}^{(\oplus)}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{j,t}),\quad j=1,\dots,J+1,$
$Y_{1,T_{0}+1:T}^{(\oplus)}(D)$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}d^{2}(\nu,Y_{1,t}(D)),\quad D\in\{N,I\}.$
@@
$Y_{2:J+1,t}^{(\bm{w})}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,t}),\quad Y_{j,1:T_{0}}^{(\bm{\lambda})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,Y_{j,t}),$
$Y_{2:J+1,1:T_{0}}^{(\bm{w},\bm{\lambda})}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,1:T_{0}}^{(\bm{\lambda})}),\quad Y_{2:J+1,T_{0}+1:T}^{(\bm{w},\oplus)}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,Y_{j,T_{0}+1:T}^{(\oplus)}).$
With $\Gamma_{\alpha,\beta}:\mathcal{M}\to\mathcal{M}$ denoting a geodesic transport map for $\alpha,\beta\in\mathcal{M}$ as introduced in Assumption S1, we define the GSDID estimator for $\tau$ as
@@
$\overline{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\Delta^{J-1}}\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}d^{2}(Y_{1,t},Y_{2:J+1,t}^{(\bm{w})}),\quad\overline{\bm{\lambda}}\in\operatorname*{arg\,min}_{\bm{\lambda}\in\Delta^{T_{0}-1}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,1:T_{0}}^{(\bm{\lambda})}).$
@@
which is the SDID estimator of Arkhangelsky et al. (2021). Thus, the proposed estimator is a natural extension of the established case with Euclidean outcomes. We discuss extensions to cases with multiple treated units or where the interest lies in the causal effects at multiple post-treatment time points in Remark S4 below.
Remark S2 (Connection with geodesic synthetic control method and geodesic difference-in-differences).
To discuss the relation of the proposed GSDID estimator with the GSC estimator introduced above and the geodesic DID (GDID) estimator introduced in Zhou et al. (2025), we note that GSC, GDID, and GSDID estimators can be represented as geodesic transports based on weighted averages of outcomes $\delta_{j,k}^{(\mathrm{M})}$, $k=1,2,3$, using appropriate weights $w_{j}^{(\mathrm{M})}$, $\mathrm{M}\in\{\mathrm{GSC},\mathrm{GDID},\mathrm{GSDID}\}$, i.e.,
@@
$Y_{1,T_{0}+1:T}^{(\mathrm{M})}=\Gamma_{\delta_{2:J+1,1:T_{0}}^{(\mathrm{M})},\delta_{2:J+1,T_{0}+1:T}^{(\mathrm{M})}}\left(\delta_{1,1:T_{0}}^{(\mathrm{M})}\right),\quad\delta_{2:J+1,1:T_{0}}^{(\mathrm{M})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}^{(\mathrm{M})}d^{2}(\nu,\delta_{j,1:T_{0}}^{(\mathrm{M})}),$
$\delta_{2:J+1,T_{0}+1:T}^{(\mathrm{M})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}^{(\mathrm{M})}d^{2}(\nu,\delta_{j,T_{0}+1:T}^{(\mathrm{M})}),\quad\delta_{1,1:T_{0}}^{(\mathrm{M})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}^{(\mathrm{M})}d^{2}(\nu,\overline{\delta}_{j,1:T_{0}}^{(\mathrm{M})}),$
@@
$\{w_{j}^{(\mathrm{GSC})},\delta_{j,1:T_{0}}^{(\mathrm{GSC})},\delta_{j,T_{0}+1:T}^{(\mathrm{GSC})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSC})}\}$ $=\{\overline{w}_{j},Y_{j,1:T_{0}}^{(\oplus)},Y_{j,T_{0}+1:T}^{(\oplus)},Y_{j,1:T_{0}}^{(\oplus)}\},$
$\{w_{j}^{(\mathrm{GDID})},\delta_{j,1:T_{0}}^{(\mathrm{GDID})},\delta_{j,T_{0}+1:T}^{(\mathrm{GDID})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GDID})}\}$ $=\{\frac{1}{J},Y_{j,1:T_{0}}^{(\oplus)},Y_{j,T_{0}+1:T}^{(\oplus)},Y_{1,1:T_{0}}^{(\oplus)}\},$
$\{w_{j}^{(\mathrm{GSDID})},\delta_{j,1:T_{0}}^{(\mathrm{GSDID})},\delta_{j,T_{0}+1:T}^{(\mathrm{GSDID})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSDID})}\}$ $=\{\overline{w}_{j},Y_{j,1:T_{0}}^{(\overline{\bm{\lambda}})},Y_{j,T_{0}+1:T}^{(\oplus)},Y_{1,1:T_{0}}^{(\overline{\bm{\lambda}})}\}.$
@@
$Y_{1,T_{0}+1:T}^{(\mathrm{GSC})}$ $=\Gamma_{Y_{2:J+1,1:T_{0}}^{(\overline{\bm{w}},\oplus)},Y_{2:J+1,T_{0}+1}^{(\overline{\bm{w}})}}\left(Y_{2:J+1,1:T_{0}}^{(\overline{\bm{w}},\oplus)}\right)=Y_{2:J+1,T_{0}+1}^{(\overline{\bm{w}})},$ (S1)
$Y_{1,T_{0}+1:T}^{(\mathrm{GDID})}$ $=\Gamma_{Y_{2:J+1,1}^{(\oplus)},Y_{2:J+1,2}^{(\oplus)}}\left(Y_{1,1}\right),$ (S2)
@@
$Y_{2:J+1,t}^{(\oplus)}$ $=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{J}\sum_{j=2}^{J+1}d^{2}(\nu,Y_{j,t}),\ t=1,2.$
The resulting estimators (S1) and (S2) coincide with the GSC and GDID estimators, respectively. Under Assumption S3 below, we have $Y_{1,T_{0}+1}^{(\mathrm{GSC})}=Y_{1,T_{0}+1}(N)$. Comparing $\{\delta_{j,1:T_{0}}^{(\mathrm{GSC})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSC})}\}$ and $\{\delta_{j,1:T_{0}}^{(\mathrm{GSDID})},\overline{\delta}_{j,1:T_{0}}^{(\mathrm{GSDID})}\}$, the GSDID can be interpreted as attenuating the bias of the GSC estimator by introducing additional flexibility through the weights $\overline{\bm{\lambda}}$ that are not present in the GSC. Under Assumption S4 below, we have $Y_{1,2}^{(\mathrm{GDID})}=Y_{1,2}(N)$. Comparing the weights $w_{j}^{(\mathrm{GDID})}$ and $w_{j}^{(\mathrm{GSDID})}$, GSDID can be seen as introducing additional flexibility to estimate the target $Y_{1,T_{0}+1:T}^{(\oplus)}(N)$ through the weights $\overline{\bm{w}}$, which may attenuate bias when the parallelity assumption does not hold. The proposed GSDID thus emerges as a method that combines favorable properties of both GSC and GDID.
S.2.2 Causal model for geodesic synthetic difference-in-differences
@@
$Y_{j,t}(N)=g(U_{j},V_{t}),\quad t=1,\ldots,T,\ j=1,\ldots,J+1,$ (S3)
@@
$g^{(\oplus)}(U_{j},V_{T_{0}+1:T})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}\!\!\!d^{2}(\nu,g(U_{j},V_{t})).$
@@
$g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,g(U_{j},V_{t})),$
$g^{(\bm{w})}(U_{2:J+1},V_{t})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g(U_{j},V_{t})),$
$g^{(\bm{w},\oplus)}(U_{2:J+1},V_{T_{0}+1:T})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g^{(\oplus)}(U_{j},V_{T_{0}+1:T})),$
$g^{(\bm{w},\bm{\lambda})}(U_{2:J+1},V_{1:T_{0}})=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})).$
@@
Assumption S2.
For any $\bm{w}\in\Delta^{J-1}$ and $\bm{\lambda}\in\Delta^{T_{0}-1}$, weighted Fréchet means $g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}})$, $g^{(\oplus)}(U_{j},V_{T_{0}+1:T})$, $g^{(\bm{w})}(U_{2:J+1},V_{j})$, $g^{(\bm{w},\oplus)}(U_{2:J+1},V_{T_{0}+1:T})$, and $g^{(\bm{w},\bm{\lambda})}(U_{2:J+1},V_{1:T_{0}})$ exist and are unique.
@@
$\mathbb{W}_{t}^{*}=\left\{\bm{w}^{*}\in\Delta^{J-1}:d(g(U_{1},V_{t}),g^{(\bm{w}^{*})}(U_{2:J+1},V_{t}))=\inf_{\bm{w}\in\Delta^{J-1}}d(g(U_{1},V_{t}),g^{(\bm{w})}(U_{2:J+1},V_{t}))\right\},$
and
$\Lambda_{j}^{*}=\Big\{\bm{\lambda}^{*}\in\Delta^{T_{0}-1}:$ $d(g^{(\oplus)}(U_{j},V_{T_{0}+1:T}),g^{(\bm{\lambda}^{*})}(U_{j},V_{1:T_{0}}))$
$=\inf_{\bm{\lambda}\in\Delta^{T_{0}-1}}d(g^{(\oplus)}(U_{j},V_{T_{0}+1:T}),g^{(\bm{\lambda})}(U_{j},V_{1:T_{0}}))\Big\}.$
Assumption S3 (Synthetic control).
@@
    $\mathbb{W}_{\bm{\lambda}}^{*}$ $=\left\{\bm{w}\in\Delta^{J-1}:d(g^{(\bm{\lambda})}(U_{1},V_{1:T_{0}}),g^{(\bm{w},\bm{\lambda})}(U_{2:J+1},V_{1:T_{0}}))=0\right\},$
    $\mathbb{W}_{\oplus}^{*}$ $=\left\{\bm{w}\in\Delta^{J-1}:d(g^{(\oplus)}(U_{1},V_{T_{0}+1:T}),g^{(\bm{w},\oplus)}(U_{2:J+1},V_{T_{0}+1:T}))=0\right\}.$
Assumption S4 (Parallel trend).
@@
Assumption S3 is for the GSC. Assumption S3(ii) implies that the optimal weights $\bm{w}$ such that $Y_{2:J+1,t}^{(\bm{w})}=Y_{1,t}(N)$ at each time point $t=1,\dots,T_{0}$ are still optimal in the sense that $Y_{1,1:T_{0}}^{(\bm{\lambda})}(N)$ and $Y_{1,T_{0}+1:T}^{(\oplus)}(N)$ respectively coincide with $Y_{2:J+1,1:T_{0}}^{(\bm{w},\bm{\lambda})}(N)$ and $Y_{2:J+1,T_{0}+1:T}^{(\bm{w},\oplus)}(N)$ as long as $\bm{\lambda}$ belongs to the set of optimal weights $\bigcap_{j=2}^{J+1}\Lambda_{j}^{*}$ such that for $j=2,\dots,J+1$, $Y_{j,1:T_{0}}^{(\bm{\lambda})}$ is the closest to $Y_{j,T_{0}+1:T}^{(\oplus)}(N)$ with respect to the metric $d$ among the Fréchet means $Y_{j,1:T_{0}}^{(\bm{\lambda}_{0})}$ with $\bm{\lambda}_{0}\in\Delta^{T_{0}-1}$. Figure S2 illustrates the GSDID under Assumption S3.
Figure S2: Illustration of geodesic synthetic difference-in-differences under Assumption S3. The circles symbolize objects in the geodesic metric space $\mathcal{M}$. Under Assumption S3, $Y_{2:4,1:3}^{(\overline{\bm{w}},\overline{\bm{\lambda}})}=Y_{1,1:3}^{(\overline{\bm{\lambda}})}$ and $Y_{2:4,4:6}^{(\overline{\bm{w}},\oplus)}=Y_{1,4:6}^{(\oplus)}$. The solid red arrow is the geodesic from $Y_{1,1:3}^{(\overline{\bm{\lambda}})}$ to $Y_{1,4:6}^{(\oplus)}(N)$, representing the bias correction for geodesic synthetic controls as effected by geodesic synthetic difference-in-differences.
Assumption S4(ii) corresponds to the parallel trend assumption for the GDID. We refer to Zhou et al. (2025) for more details about its construction and theoretical properties. For Euclidean outcomes, Assumption S4(ii) reduces to
@@
which corresponds to the standard assumption of parallel trends in the canonical DID framework (Zhou et al., 2025). Figure S3 illustrates the GSDID under Assumption S4.
Figure S3: Illustration of GSDID under Assumption S4. The circles symbolize objects in the geodesic metric space $\mathcal{M}$. Arrows depict geodesics connecting them. The dashed red arrow from $Y_{1,1:3}^{(\overline{\bm{\lambda}})}$ to $Y_{1,4:6}^{(\oplus)}(N)$ represents the trend between pre-treatment and post-treatment for unit 1. The solid red arrow represents the parallel trend. The dashed red arrow from $Y_{2:4,4:6}^{(\overline{\bm{w}},\oplus)}$ to $Y_{1,4:6}^{(\oplus)}(N)$ represents the bias correction of the GSC affected by the GSDID.
For the space of networks (Example 1), Wasserstein space (Example 4), the space of functional data (Example 5), and the space of SPD matrices (Example 3) with either the Frobenius metric, the power metric family, the Log-Euclidean metric or the Log-Cholesky metric, one can verify Assumptions S2, S3, and S4(i). Indeed, letting $g(U_{j},V_{t})=\gamma_{U_{j},V_{t}}(\alpha)$, $\alpha\in(0,1)$,
@@
$U_{2:J+1}^{(\bm{w})}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{j=2}^{J+1}w_{j}d^{2}(\nu,U_{j}),$
$V_{T_{0}+1:T}^{(\oplus)}=\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}\!\!\!d^{2}(\nu,V_{t}),\ V_{1:T_{0}}^{(\bm{\lambda})}\in\operatorname*{arg\,min}_{\nu\in\mathcal{M}}\sum_{t=1}^{T_{0}}\lambda_{t}d^{2}(\nu,V_{t}).$
For the space of compositional data (Example 2), consider the causal model $g(U_{j},V_{t})=\mathrm{Exp}_{V_{t}}(\mathrm{Log}_{U_{1}}(U_{j}))$. If we additionally assume that